OpenStudy (anonymous):

What is the solution to this equation? 16x - 10x + 9 = 99 - 30 A. x = 10 B. x = 11 C. x = 13 D. x = 12

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

16x- 10x= 6x....99-9=90, 90/6= 15!! so its non of the above

OpenStudy (anonymous):

It's definitely A because if you plug it in, it works out correctly: 16(10)- 10(10)+9 = 99-30

OpenStudy (anonymous):

whoops didnt see that -30, yup @smartkidd18 is right

OpenStudy (anonymous):

how are yall doing it lol

OpenStudy (anonymous):

I just used a calculator. I replaced the x for each of the answer choices to see which would fit. Its the Process of elimination basically. You could solve it mathematically but in my opinion, it takes longer.